OpenStudy (a1234):

Find the values of the trig functions if sin t = (sqrt13)/7 and the terminal point of t is in Quadrant 2. My answers are... cot t = -(6(sqrt13))/13 tan t = -(sqrt13)/6 cos t = -6/7

OpenStudy (jdoe0001):

cot t = -(6(sqrt13))/13 \(\large \checkmark\) tan t = -(sqrt13)/6 \(\large \checkmark\) cos t = -6/7 \(\large \checkmark\)

OpenStudy (a1234):

Thanks!